OpenStudy (beccab003):

Choose the correct equation of a line, in point-slope form, given the point (-1,3) and a slope, m, of 1/2. 1) y+1=1/2(x-3) 2) y-3=1/2(x+1) 3) y+3=1/2(x-1) 4) y-1=1/2(x-3)

OpenStudy (aravindg):

Looks familiar? did u try to look at my answer?

OpenStudy (beccab003):

Yes, but it didn't make sense and I asked follow up questions but you didn't respond. What is the answer?

OpenStudy (beccab003):

I think its 3

OpenStudy (anonymous):

Why did you pick 3?

OpenStudy (aravindg):

You have y-y1 in equation here y1=3 So it should be y-3 instead of y+3

OpenStudy (beccab003):

Cause the 3 is positive and the 1 is negative

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\] point-slope form

OpenStudy (aravindg):

3 is positive but we have y-y1 y-3

OpenStudy (anonymous):

^

OpenStudy (beccab003):

So is it number 2

OpenStudy (anonymous):

Yes

OpenStudy (beccab003):

But how come its x+1 when 1 is negative?

OpenStudy (anonymous):

Look at the formula